compare the graph of the function with the graph of f(x)=x^2.
g(x)=x^2-3
width :
opens up or down:
vertex :
thanks!

Answer by jim_thompson5910(35256)

Take x^2 and shift it down 3 units


The "width" will stay the same since we're not compressing the graph in any way.


The vertex for x^2 is (0,0) and it shifts down to (0,-3) for y = x^2 - 3


Both graphs open upwards


Graph of y = x^2 and y = x^2 - 3